a square piece of paper is folded in half. The perimeter of the resulting rectangle is 24cm. Find the area of the origional paper
Answer by mananth(16946)

Let the side of square be x
When the paper is folded the resultant rectangle will have length same as side of the square and the width 1/2 of the square side
Perimeter of rectangle = x+x+x/2+x/2 = 3x
3x=24
x=8
The area of originalaper is 8*8 = 64 square cm